Circular No. 5751

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993J IN NGC 3031 R. Fesen and J. Thorstensen, Dartmouth College, report that spectra

(range 440-800 nm) of SN 1993J taken Mar. 31.1 and Apr. 1.1 UT using the MDM Observatory 2.4-m telescope show faint, broad emission features with peaks near 495, 586, and 656 nm. If the 656-nm feature is due to Hα, its FWZI implies a Doppler width of ± 10 000 km/s; its equivalent width increased from 0.7 nm on Mar. 31 to 1.0 nm on Apr. 1. The 586-nm feature extends to 565 nm, but its red edge was indistinct. If interpreted as He I 587.6 nm, the blue extent corresponds to a velocity of 12 000 km/s.

S. Benetti, G. Contarini, R. Gratton and M. Turatto, Asiago Astrophysical Observatory, communicate: “We have taken high-resolution (0.035 nm) spectra of SN 1993J with the échelle spectrograph at Cima Ekar on Mar. 31.0 UT. The narrow Hα emission at 656.14 nm (corresponding to a heliocentric velocity of –82 km/s) is resolved (FWHM 0.45 nm, or 200 km/s), and the EW is 0.034 nm, in agreement with that on IAUC 5736. No defined structure of the line is shown. We also observed weak emissions due to He II at 468.41 nm (EW 0.008 nm), [Fe X] at 637.22 nm (FWHM 0.10 nm, EW = 0.009 nm), and [Fe XIV] at 530.06 nm (FWHM 0.09 nm, EW = 0.003 nm). Various absorption components of the Na I D lines were observed (at –115 km/s, EW(D_{1}) = 0.0164 nm, EW(D_{2}) = 0.0127 nm; at +133 km/s, EW(D_{1}) = 0.0396 nm, EW(D_{2}) = 0.0331 nm), in good agreement with IUE observations of Mg I on IAUC 5738. Other blended components are also present. The Na I column densities for the clouds responsible for the two measured components are log n(Na I) = 12.48 and 13.02, respectively, using the doublet ratio method and gaussian line profiles. This suggests a reddening of E(B–V) between 0.1 and 0.4 (Gratton and Sneden 1987, A.Ap. 204, 193).”

G. G. Pooley and D. A. Green, Mullard Radio Astronomy Observatory, Cambridge, write: “Using observations made with the Ryle telescope during Apr. 5.631-5.898 UT, we have detected radio emission from SN 1993J at a frequency of 15 GHz. The flux density was ∼ 3 mJy, with strong evidence for increasing flux density during the observations. No emission was detected during observations at similar hour angles the previous day. The position of the detected emission is within 1′′ of the optical position reported on IAUC 5737.”

1993 April 6 Daniel W. E. Green